When two bodies move uniformly towards each other, the distance decreases by 6m/s. If both bodies move 8n the same direction with same speeds the distance b/w them increases by 4m/s. Then the speeds of two bodies are

Grade:12

1 Answers

TECH HACK
6 years ago
Let the velocity of first one be A and that of second be B .Then in first case going in opposite directions therefore there relative velocity will be A+B=6And in second case going in same direction therefore there relative velocity will be A-B=4m/s.Solving the equation we get A =5m/s and B=1m/s . That`s the velocity of the body.Hope it helps.